OpenStudy (impracticaljoker):

8th grade math, I'd appreciate some help. ;)

rebeccaxhawaii (rebeccaxhawaii):

whats up

OpenStudy (impracticaljoker):

Jordan wants to solve the following system using the elimination method: 2x + 3y = 10 x + y = 7 What number should the equation x + y = 7 be multiplied by to eliminate y? (4 points) –2 2 –3 3

OpenStudy (4everaddicted2anime):

In order to eliminate y, both y's have to have the same coefficient. You have y and then you have 3y. What do you have to multiply y by to get 3y?

OpenStudy (phi):

this is a lousy question. You could multiply the bottom equation by 3 and then subtract, i.e. do top equation - bottom equation or, you could multiply the bottom equation by -3 and then add: top equation plus bottom equation. I am *guessing* they assume you add the two equations, but who knows? I would complain to the teacher about this.

Directrix (directrix):

2x + 3y = 10 =========> 2x + 3y = 10 x + y = 7 --> by -3 ====> -3x - 3y = -21 ----------------- Add the two equations and see what happens to the y-term. @ImpracticalJoker Note: You could also multiply by 3 and then subtract the second equation from the first. Both options are listed. That is what makes this a lousy question.